OpenStudy (anonymous):

Find the commission for a rate of 5.5% for $1,400 in sales.

OpenStudy (anonymous):

You can solve it like this: 5.5% = 5.5/100 = 0.055 (in decimal format). Then multiply: 0.055 * $1,400 = $77 Therefore, the commission is $77 Hope this helps.